OpenStudy (anonymous):

find the derivative of derivative of (x - 1)^2 * (x - 2)^3

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{d }{dx} \left[ f(x) {\color{white}{\Huge |}}\cdot g(x)\right] =f(x)g'(x)+f'(x)g(x) }\)

OpenStudy (solomonzelman):

First, you would want to find the derivative of each function, and then apply the product rule (that I posted above)

OpenStudy (anonymous):

@SolomonZelman would I use the chain rule on (x-1)^2 ??

OpenStudy (solomonzelman):

technically you are, but the derivative of x-1 and of x-2 is 1, so to write out that you are multiplying times 1, is a waste of time.